OpenStudy (anonymous):

What is the slope of the line passing through (1, 2) and (3, 8)?

OpenStudy (anonymous):

The formula for finding slope is y2-y1/x2-x1 Just plug the points into this formula (1, 2) (3, 8) (8-2)/(3-1)= 6/2 = 3 The slope would be 3, therefore the answer is C